OpenStudy (anonymous):

A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet. a. Find the length of the side of the lot opposite the 60° angle. b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (anonymous):

i really dont know what im doing wrong...

OpenStudy (anonymous):

 (a) If the "1" side is 41, then the side opposite 60-degrees is sqrt(3). So L = \[41\sqrt{3}\] or 71.0141 feet (b) If the hypotenuse is the "2" side then simply double the "1" side. So L = 41*2 = 82 feet (c) sin(30) = (41/82) = 0.5000 cos(30) = (71.0141/82) = 0.8660 tan(30) = (41/71.0141) = 0.5774